OpenStudy (anonymous):

Given f '(x) = (x - 4)(4 - 2x), find the x-coordinate for the relative maximum on the graph of f(x) A)4 B)3 c)2 D)None of these

OpenStudy (whpalmer4):

What will be true about the value of \(f'(x)\) at a relative maximum?

OpenStudy (anonymous):

It will be (3,2)?

OpenStudy (whpalmer4):

Is that an answer to my question, or to the problem?

OpenStudy (anonymous):

To your question

OpenStudy (whpalmer4):

so you're saying that for any function, the value of \(f'(x)\) at a relative maximum is (3,2)? :-)

OpenStudy (whpalmer4):

I'm asking a general question here...what do we know about the first derivative at a local maximum?

OpenStudy (anonymous):

Oh I thought your question was about my specific question

OpenStudy (whpalmer4):

we will use it to get an answer to your specific question.

OpenStudy (anonymous):

Well it changes sign at the relative maximum, right?

OpenStudy (whpalmer4):

the first derivative is the instantaneous slope of the tangent to the function. so at a local maximum or a local minimum, it will be equal to 0.

OpenStudy (anonymous):

So how do I apply it to this question?

OpenStudy (whpalmer4):

where are the points where \(f'(x) = (x-4)(4-2x) = 0\)?

OpenStudy (anonymous):

Oh... Yeah that makes sense, so the answer would be 2 right? Thank you a lot for your help!

OpenStudy (whpalmer4):

not so fast... is that the only place where \(f'(x) = (x-4)(4-2x) = 0\)?

OpenStudy (anonymous):

No, there are two points where it equals 0

OpenStudy (whpalmer4):

by using the zero product rule, \[(x-4)(4-2x) = 0\]has solutions at \[(x-4) = 0 \rightarrow x = 4\]\[(4-2x) = 0\rightarrow x = 2\] We just know (so far) that \(x = 4\) and \(x=2\) are either local maximum or local minimum points on the graph of \(f(x\).

OpenStudy (whpalmer4):

How do you propose to figure out which one is the local maximum?

OpenStudy (anonymous):

I don't really know. What do I do? :(

OpenStudy (anonymous):

Whapalmer!

OpenStudy (whpalmer4):

do you know how to integrate a polynomial?

OpenStudy (whpalmer4):

if you do, you could integrate \(f'(x)\) and evaluate the result (which would be \(f(x) + C\) where C is an arbitrary constant) at both values of \(x\) and see which one is the local maximum and which is the local minimum. That would one approach. Another would be to look at the value of \(f'(x)\) between the two inflection points (points where \(f'(x) = 0\)). If that value is positive, that means the slope of the tangent to \(f(x)\) is also positive. What would imply about which point was the local maximum?

OpenStudy (anonymous):

is it 4?

OpenStudy (whpalmer4):

Well, I'll tell you if it is or not, if you tell me why you think 4 is the answer...

OpenStudy (whpalmer4):

Let's consider the value of \(f'(1)\). \(f'(x) = (x-4)(4-2x)\) so \(f'(1) = (1-4)(4-2(1)) = -6\) which means that the function is decreasing at \(x = 1\). Now let's consider \(f'(3) = (3-4)(4-2(3)) = 2\) which means the function is increasing at \(x = 3\). And finally \(f'(5) = (5-4)(4-2(5)) = -6\) so the function is again decreasing at \(x=5\). \[\begin{array}{cc} x &f(x)\text{ is increasing/decreasing}\\ \hline\\1 & \text{decreasing}\\ 2&\text{inflection point}\\ 3&\text{increasing}\\ 4&\text{inflection point}\\ 5&\text{decreasing} \end{array} \] Hopefully, from that table you can visualize that \(f(2\) is a local minimum, and \(f(4)\) is a local maximum.